OpenStudy (anonymous):

The graph of g(x) is f(x) translated to the left 8 units and up 2 units. What is the function rule for g(x) given f(x)=x^2

OpenStudy (anonymous):

@jdoe0001 help

OpenStudy (anonymous):

Are you able to help @pgpilot326

OpenStudy (anonymous):

http://www.mathsisfun.com/sets/function-transformations.html have a look at this site and see if that gets you where you need to be. if you need help after, i'll be here

OpenStudy (anonymous):

no i know all of that im just having trouble solving this problem.

OpenStudy (anonymous):

so what is h? what is k?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

g(x) = f(x-h) + k

OpenStudy (anonymous):

44 JjJustin Medals 0 no i know all of that im just having trouble solving this problem. apparently you don't know all of that then.

OpenStudy (shamil98):

g(x)=f(x) = (x^2+8) +2

OpenStudy (shamil98):

the + 8 = left 8 units , this is what you call a phase shift .

OpenStudy (anonymous):

i have 1 medal. and stop stalking my profile queer

OpenStudy (anonymous):

ok thanks shamil. that means i did one problem wrong xD now i have to go back and change some work

OpenStudy (anonymous):

yo... it's called copy and paste to demostrate how you don't know all that. i guess you're not here to learn because you "know all that."